# Universal Efimov Scaling in the Rabi-Coupled Few-Body Spectrum

### v1

@@ Abstract
A key property of all these manifestations of Efimov physics is the three-body parameter, a short-distance scale which sets the deepest bound trimer and hence uniquely determines the entire Efimov spectrum due to the trimers’ discrete scaling symmetry Braaten and Hammer (2006); Naidon and Endo (2017). Specifically, by knowing the scattering length $a_{-}<0$ at which the deepest trimer unbinds, we also know this parameter for all excited Efimov trimers via the relation $a_{-}^{(n)}=\lambda^{n}a_{-}$, where $n=1,2,\dots$ and $\lambda$ is a universal scaling factor. Similarly, at the two-body resonance $a\to\pm\infty$, the trimer energies satisfy $E_{0}^{(n)}=\lambda^{-2n}E_{0}$, where $E_{0}$ is the deepest trimer energy. Remarkably, it has also been shown in the cold-atom context that there is a fixed relationship between the three-body parameter and the underlying short-distance physics, parameterized by the van der Waals range $R_{\mathrm{vdW}}$, e.g., for the canonical case of three identical bosons, $a_{-}\simeq-9.5R_{\mathrm{vdW}}$ Berninger et al. (2011). This so-called van der Waals universality even holds in the presence of dipolar interactions Oi et al. (2024). Therefore, an important question is whether this relation can be modified, ultimately yielding a tunable three-body parameter.
Figure 1: (a) Schematic of the Rabi-coupled three-body problem. The Rabi-dressed atom (purple) consists of a superposition of a spin-$\uparrow$ state (red), which interacts with the heavy bosons (green), and a non-interacting spin-$\downarrow$ state (blue). (b) The Efimov spectrum with (solid lines) and without (dashed) the Rabi drive for the mass ratio $133/6$. The axes represent the original parameters (left and top) and the effective parameters in the Rabi-driven case (bottom and right), corresponding to the effective scattering length $a_{\rm eff}\propto(1/a-1/a_{\rm c})^{-1}$ and the energy $\delta E=E-{\epsilon_{-}}$ measured from the Rabi-shifted continuum. The uncoupled Efimov spectrum is centered around $(1/a=0,E=0)$, while the coupled spectrum is shifted (arrow) to be centered around $(1/a=1/a_{\text{c}},E={\epsilon_{-}})$. The two-body bound state (purple), trimers (blue) and tetramers (pink) are visible in both spectra. The Rabi-coupled Efimov spectrum is calculated for $ma_{-}^{2}\Omega_{0}=2$ and $\Delta_{0}/\Omega_{0}=-1$.
@@
As shown in Fig. 1(b), we find that the Rabi drive shifts the spectrum of Efimov trimers, and associated four-body bound states (tetramers), such that it lies around the new two-body resonance Mulkerin et al. (2024); Ble for the Rabi-dressed impurity. In particular, we find that the infinite tower of trimer states, along with the usual discrete scaling symmetry, is preserved by the Rabi coupling, unlike the related scenario of spin-orbit coupling Shi et al. (2014, 2015); Guan and Blume (2018), where there is instead a modified scaling law in a higher dimensional space Guan and Blume (2018). We also show that the whole spectrum can be conveniently explored by fixing the scattering length $a$ and instead varying the detuning of the Rabi drive from the atomic transition. Crucially, we find that the three-body parameter is universally shifted by the Rabi drive, with a log-periodic dependence on the strength of the Rabi coupling.
@@
where $\mathbf{k}$ is the particle’s momentum, $\epsilon_{\mathbf{k}}^{b}=|\mathbf{k}|^{2}/2M\equiv k^{2}/2M$ is the dispersion for the identical bosons with mass $M$, and $\epsilon_{\mathbf{k}}^{c}=k^{2}/2m$ is the corresponding dispersion for the light impurity with mass $m$ and spin $\sigma\in\{\uparrow,\downarrow\}$. We work in units such that $\hbar$ and the system volume are both set to 1. We focus on the experimentally accessible case of a 133Cs-6Li mixture Tung et al. (2014); Pires et al. (2014); Lippi et al. (2024), where the mass ratio $M/m=133/6$, resulting in an Efimov scaling factor of $\lambda=4.87$ Braaten and Hammer (2006); Naidon and Endo (2017) when the boson-boson interactions are negligible. For simplicity, we also assume that the spin-$\downarrow$ state is completely non-interacting, but this condition can straightforwardly be relaxed Adlong et al. (2020); Hu and Liu (2023); Ble .
@@
Here, $\Omega_{0}$ is the Rabi coupling and $\Delta_{0}$ is the detuning from the bare $\uparrow$-$\downarrow$ transition. Solving the single-particle problem yields two Rabi-split quasiparticle branches with dispersions $E_{\mathbf{k}}^{\pm}=\epsilon_{\mathbf{k}}^{c}+{\epsilon_{\pm}}$, where ${\epsilon_{\pm}}=\left(\Delta_{0}\pm\sqrt{\Omega_{0}^{2}+\Delta_{0}^{2}}\right)/2$, for the upper and lower branches, respectively. Their respective (real) spin-$\uparrow$ amplitudes $v$ and $u$ satisfy $u^{2}=\left(1+\Delta_{0}/\sqrt{\Omega_{0}^{2}+\Delta_{0}^{2}}\right)/2$ and $u^{2}+v^{2}=1$.
@@
$\hat{H}_{\uparrow}=\sum_{\mathbf{k}}\epsilon_{\mathbf{k}}^{{\vphantom{\dagger}}d}\hat{d}^{\dagger}_{\mathbf{k}}\hat{d}^{\vphantom{\dagger}}_{\mathbf{k}}+g\sum_{\mathbf{k}\mathbf{q}}\left(\hat{d}^{\dagger}_{\mathbf{q}}\hat{c}^{\vphantom{\dagger}}_{\mathbf{q}-\mathbf{k}\uparrow}\hat{b}^{\vphantom{\dagger}}_{\mathbf{k}}+\hat{b}^{\dagger}_{\mathbf{k}}\hat{c}^{\dagger}_{\mathbf{q}-\mathbf{k}\uparrow}\hat{d}^{\vphantom{\dagger}}_{\mathbf{q}}\right),$ (4)
corresponding to a two-channel model Timmermans et al. (1999) in which two atoms can form a closed channel dimer $\hat{d}^{\dagger}_{\mathbf{k}}$ with kinetic energy $\epsilon_{\mathbf{k}}^{d}=k^{2}/2(m+M)$ and detuning $\nu_{0}$. This model has the exact low-energy scattering amplitude $f(k)=-1/(a^{-1}+R^{*}k^{2}+ik)$, where the scattering length $a$ and range parameter $R^{*}$ replace the bare model parameters via $m_{r}/2\pi a=-\nu_{0}/g^{2}+\sum_{\mathbf{k}}1/(\epsilon^{b}_{\mathbf{k}}+\epsilon^{c}_{\mathbf{k}})$ and $R^{*}=\pi/m_{r}^{2}g^{2}$, where $m_{r}=mM/(m+M)$ is the reduced mass. The introduction of the closed channel is useful, as it provides a UV cutoff that regulates the Efimov physics, thus introducing an effective three-body parameter. Indeed, for the specific case of a 133Cs-6Li mixture, the deepest bound Efimov trimer unbinds at the scattering length $a_{-}=-1.04R^{*}$ and has energy $E_{0}=-0.242mR^{*2}$ at unitarity. However, as explicitly shown in the Supplemental Material (SM) SM , our results in the following do not sensitively depend on the choice of UV cutoff  [^1] Our model aims to capture the essential physics of Rabi-coupled mass-imbalanced mixtures. For direct comparison with Cs-Li experiments, one may need to employ more complicated models in order to capture overlapping Fano-Feshbach resonances Repp et al. (2013); Tung et al. (2013), such as models involving three interaction channels Johansen et al. (2017); Yudkin and Khaykovich (2021); Li et al. (2022)..
@@
where $a_{\rm{eff}}$ diverges at $a=a_{\rm{c}}$, as expected. Note that this differs from the resonances induced by sinusoidally driving the scattering length Smith (2015), since there is no imaginary part to $a_{\rm{eff}}$.
@@
where $E^{\pm}_{\mathbf{k}_{1}\mathbf{k}_{2}}=\epsilon_{\mathbf{k}_{1}}^{b}+\epsilon_{\mathbf{k}_{2}}^{b}+E_{-\mathbf{k}_{1}-\mathbf{k}_{2}}^{\pm}$ and $\gamma_{\mathbf{k}}$ is a three-body vertex function. Similar to the standard three-body problem, we find that the solutions of Eq. (8) yield a spectrum with infinitely many three-body bound states. As shown in Fig. 1(b), these are centered around the new resonance and become unbound at the shifted continuum. In order to directly compare with the usual Efimov scenario, we plot the Rabi-coupled spectrum in terms of the effective scattering length, Eq. (7), and measure energy with respect to the shifted continuum, $\delta E=E-{\epsilon_{-}}$. We see that the spectrum resembles the undriven case, with some modifications for the tightest bound states. In particular, we find that the excited states satisfy the usual scaling relations for the critical scattering lengths at which the Efimov trimers unbind, $a_{\mathrm{eff},-}^{(n)}=\lambda a_{\mathrm{eff},-}^{(n-1)}$, and for their energies at unitarity, $\delta E^{(n)}=\lambda^{-2}\delta E^{(n-1)}$.
Figure 2: (a) Tuning the three-body parameter via a Rabi drive. We show the energies (solid lines) of the first ten excited Efimov trimers at the induced resonance ($1/a=1/a_{\rm{c}}$) and at $\Delta_{0}=0$ as a function of Rabi-coupling strength, ignoring the less universal ground state. Here, we have scaled the energies $\delta E^{(n)}$ by $\lambda^{2n}$ so that the lines collapse in the asymptotic large-$n$ limit, where the gray dashed line corresponds to the behavior in the absence of Rabi coupling: $E_{0}^{(n)}\lambda^{2n}=-0.194/mR^{*2}$. The purple dashed line is the Rabi-induced universal behavior obtained from the Born-Oppenheimer (BO) approximation SM . (b) The difference $\delta V(R)$ between the coupled and uncoupled BO potentials at $01/a_{\mathrm{eff}}=$ 0 and $01/a=$ 0, respectively, as a function of the distance $R$ between the two heavy atoms. Here, we take $R^{*}=0$ for both potentials and $\Delta_{0}=0$ for the coupled potential.
@@
with $s_{0}\simeq W(1)\sqrt{M/2m}$ and $k_{\Omega}=\sqrt{2m\Omega_{0}}$. For a given Efimov trimer, the ratio of binding energies with and without the Rabi drive is given by $\exp\left(2[\delta(R\to\infty)-\delta(R_{0})]/s_{0}\right)$ where $R_{0}$ is a short-range scale SM . Applying this analysis directly yields the purple line in Fig. 2(a), which all trimers with $|E_{0}^{(n)}|\ll\Omega_{0}$ will follow. This matches our exact numerics very well, with small deviations arising from the use of the BO approximation. We have thus demonstrated that the Rabi-dressed three-body parameter is a log-periodic function of the Rabi coupling, and hence that the three-body parameter is tunable via an external drive. The range of possible values is further increased by considering also the role of the detuning SM .
@@
Discussion.— Experimentally, one can probe the Efimov spectrum under a continuous Rabi drive by starting in the non-interacting state at $\Delta_{0}/\Omega_{0}\ll-1$ and then adiabatically varying the detuning $\Delta_{0}$ at fixed coupling $\Omega_{0}$, as in Fig. 3. The Efimov states will then show up as loss resonances whenever a bound trimer or tetramer state enters the Rabi-shifted continuum, similar to how Efimov physics is typically probed using magnetic fields Kraemer et al. (2006). However, since the driving field is typically locked to an atomic frequency standard, this method of exploring Efimov states can potentially yield clearer signatures of highly excited states than in magnetic-field-based experiments. In order to address all the excited Efimov trimers up to the first in a CsLi mixture, i.e., to have $\Omega_{0}$ as large as $|E_{0}^{(1)}|$, we estimate that we need $\Omega_{0}\lesssim 2\pi\times 100\mathrm{kHz}$, which has recently been achieved in Ref. Vivanco et al. (2023). We therefore expect that the modified three-body parameter predicted here can be observed in current or near-future experiments on strongly mass-imbalanced quantum gases.
While our work has focussed on the behavior in the steady-state regime, there is also the prospect of using a strong Rabi drive to explore Efimov physics in the time domain, e.g., via Rabi oscillations. This could naturally be implemented for the Bose polaron scenario, where a small number of light impurities are immersed in a gas of heavy bosons. Thus, our few-body results also have potential ramifications for many-body systems such as Bose-Fermi mixtures.
@@ Acknowledgements.
[…]
@@
where ${\epsilon_{\pm}}=(\Delta_{0}\pm\sqrt{\Omega_{0}^{2}+\Delta_{0}^{2}})/2$, and the real amplitudes $v$ and $u$ satisfy $u^{2}=\left(1+\Delta_{0}/\sqrt{\Omega_{0}^{2}+\Delta_{0}^{2}}\right)/2$ and $u^{2}+v^{2}=1$. Finally, replacing the bare interaction parameters with the physical scattering parameters and performing the sum over momenta, we arrive at the following equation:
@@
To look at the effective scattering parameters introduced due to the Rabi-coupling, we transform from the spin-basis to the Rabi-dressed basis; specifically, we consider the lower Rabi coupled dressed basis interaction:
@@
where we now have an effective scattering length, $a_{\rm{eff}}$, and a modified range parameter $R^{*}_{\rm{eff}}$. This dressed scattering expansion is valid when we have $|\delta E|<\Omega_{0}$. These scattering parameters are both functions of the Rabi-coupling parameters and take the form:
@@
where $1/a_{\rm{c}}=v^{2}\sqrt{2m_{r}({\epsilon_{+}}-{\epsilon_{-}})}-2R^{*}{\epsilon_{-}}m_{r}$ is the critical scattering length presented in Eq. (6) in the main text. Here, we see that $a_{\rm{eff}}\to\infty$ when $1/a\to 1/a_{\rm{c}}$, as expected, and that the range parameter is modified due to the coupling. In particular, we note that even when $R^{*}\to 0$ we have $R^{*}_{\text{eff}}\neq 0$ in the presence of a Rabi-coupling. However, as we discuss below, this Rabi-induced range parameter does not set a new three-body parameter, since this is a purely low-energy property, whereas the Efimov spectrum also relies on the behavior at short distances (high energies).
@@
In the main text, we have presented our results within the two-channel model which includes a finite effective range, thus impacting the two-body physics as well as providing an ultraviolet (UV) cutoff for the Efimov spectrum. To show that our key results are model independent, here we instead present all figures using an explicit UV cut-off that only affects the three- and four-body physics Yoshida et al. (2018a, b). To this end, we take the coupling $g$ to infinity while keeping the scattering length $a$ fixed in such a way that $R^{*}=0$. In this limit, the two-body physics coincides with that of a single-channel model. To still provide a three-body parameter for the Efimov spectrum, we follow Ref. Yoshida et al. (2018a) and introduce a UV cut-off $\Lambda$ on the momentum-exchange sums on the right-hand side of the few-body equations (8) and (9) of the main text. The introduction of such a cut-off is equivalent to considering an explicit three-body repulsion Bedaque et al. (1999). To distinguish the two models, we will denote the model with a three-body cut-off as the $\Lambda$ model, as opposed to the two-channel model considered in the main text, which we term the $R^{*}$ model. The deepest bound Efimov cluster for the 133Cs -6Li mixture now enters the scattering continuum at $a_{-}=179.578/\Lambda$.
Figure S1: (a) The Efimov spectra with (solid) and without (dashed) Rabi coupling calculated within the $\Lambda$ model. The two sets of axes represent the uncoupled case (left, top) and the Rabi-driven case (bottom, right), where the latter is expressed in units of the effective scattering length (see Eq. (7) of the main text) and the shifted continuum ($\delta E=E-{\epsilon_{-}}$). The parameters are chosen to correspond to a Cs-Li mixture as in the main text, with $ma_{-}^{2}\Omega_{0}=2$ and $\Delta_{0}/\Omega_{0}=-1$. The uncoupled Efimov spectrum (dashed) is centered around $(0,0)$, while the coupled spectrum (solid) is shifted, and centered instead at $(a_{\text{crit}},{\epsilon_{-}})$ (see arrow). The two-body bound state (purple), three-body trimers (blue) and four-body tetramers (pink) are visible in both spectra. (b) The few-body Rabi-coupled Efimov spectrum as a function of detuning, calculated within the $\Lambda$ model. The parameters correspond to a Cs-Li mixture as in the main text, with $ma_{-}^{2}\Omega_{0}=1$ and $a=|a_{-}|$. Here, we have shaded the scattering continuum (green), the continuum above the dimer (purple) and the continuum above the ground state trimer (blue). The two-body bound state (purple), trimers (blue) and tetramers (pink) all enter the continuum at the lower Rabi-coupled state, with energy ${\epsilon_{-}}$ (dashed line). Inset: The Efimov spectrum centered around the critical detuning ${\Delta_{\text{c}}}$ as a function of $\delta\Delta=\Delta_{0}-{\Delta_{\text{c}}}$.

### v2

@@ Abstract
A key property of all these manifestations of Efimov physics is the three-body parameter, a short-distance scale which sets the deepest bound Efimov trimer and hence uniquely determines the entire Efimov spectrum due to the trimers’ discrete scaling symmetry Braaten and Hammer (2006); Naidon and Endo (2017). Specifically, by knowing the scattering length $a_{-}<0$ at which the deepest trimer unbinds, we also know this parameter for all excited Efimov trimers via the relation $a_{-}^{(n)}=\lambda^{n}a_{-}$, where $n=1,2,\dots$ and $\lambda$ is a universal scaling factor. Similarly, at the two-body resonance $a\to\pm\infty$, the trimer energies satisfy $E_{0}^{(n)}=\lambda^{-2n}E_{0}$, where $E_{0}$ is the deepest trimer energy. Remarkably, it has also been shown in the cold-atom context that there is a fixed relationship between the three-body parameter and the underlying short-distance physics, parameterized by the van der Waals range $R_{\mathrm{vdW}}$, e.g., for the canonical case of three identical bosons, $a_{-}\simeq-9.5R_{\mathrm{vdW}}$ Berninger et al. (2011). This so-called van der Waals universality even holds in the presence of dipolar interactions Oi et al. (2024). Therefore, an important question is whether this relation can be modified, ultimately yielding a tunable three-body parameter.
Figure 1: (a) Schematic of the Rabi-coupled three-body problem. The Rabi-dressed atom (purple) consists of a superposition of a spin-$\uparrow$ state (red), which interacts with the heavy bosons (green), and a non-interacting spin-$\downarrow$ state (blue). (b) The Efimov spectrum with (solid lines) and without (dashed) the Rabi drive for the mass ratio 133/6133/6133 / 6. The axes represent the original parameters (left and top) and the effective parameters in the Rabi-driven case (bottom and right), corresponding to the effective scattering length $a_{\rm eff}\propto(1/a-1/a_{\rm c})^{-1}$ and the energy $\delta E=E-{\epsilon_{-}}$ measured from the Rabi-shifted continuum. The uncoupled Efimov spectrum is centered around $(1/a=0,E=0)$, while the coupled spectrum is shifted (arrow) to be centered around $(1/a=1/a_{\text{c}},E={\epsilon_{-}})$. The two-body bound state (purple), trimers (blue) and tetramers (pink) are visible in both spectra. The Rabi-coupled Efimov spectrum is calculated for $ma_{-}^{2}\Omega_{0}=2$ and $\Delta_{0}/\Omega_{0}=-1$.
@@
As shown in Fig. 1(b), we find that the Rabi drive shifts the spectrum of Efimov trimers, and associated four-body bound states (tetramers), such that it lies around the new two-body resonance Mulkerin et al. (2024); Bleu et al. (2025) for the Rabi-dressed impurity. In particular, we find that the infinite tower of trimer states, along with the usual discrete scaling symmetry, is preserved by the Rabi coupling, unlike the related scenario of spin-orbit coupling Shi et al. (2014, 2015); Guan and Blume (2018), where there is instead a modified scaling law in a higher dimensional space Guan and Blume (2018). We also show that the whole spectrum can be conveniently explored by fixing the scattering length $a$ and instead varying the detuning of the Rabi drive from the atomic transition. Crucially, we find that the three-body parameter is universally shifted by the Rabi drive, with a log-periodic dependence on the strength of the Rabi coupling.
@@
where $\hat{b}^{\dagger}_{\mathbf{k}}$ ($\hat{c}^{\dagger}_{\mathbf{k}}$) creates a heavy (light) particle with momentum $\mathbf{k}$, $\epsilon_{\mathbf{k}}^{b}=|\mathbf{k}|^{2}/2M\equiv k^{2}/2M$ is the dispersion for the identical bosons with mass $M$, and $\epsilon_{\mathbf{k}}^{c}=k^{2}/2m$ is the corresponding dispersion for the light impurity with mass $m$ and spin $\sigma\in\{\uparrow,\downarrow\}$. We work in units such that $\hbar$ and the system volume are both set to 1. We focus on the experimentally accessible case of a 133Cs-6Li mixture Tung et al. (2014); Pires et al. (2014); Lippi et al. (2024), where the mass ratio $M/m=133/6$, resulting in an Efimov scaling factor of $\lambda=4.87$ Braaten and Hammer (2006); Naidon and Endo (2017) when the boson-boson interactions are negligible. For simplicity, we also assume that the spin-$\downarrow$ state is completely non-interacting, but this condition can straightforwardly be relaxed Adlong et al. (2020); Hu and Liu (2023); Bleu et al. (2025); Mulkerin et al. (2025).
@@
Here, $\Omega_{0}$ is the Rabi coupling and $\Delta_{0}$ is the detuning from the bare $\uparrow$-$\downarrow$ transition. Solving the single-particle problem yields two Rabi-split quasiparticle branches with dispersions $E_{\mathbf{k}}^{\pm}=\epsilon_{\mathbf{k}}^{c}+{\epsilon_{\pm}}$, where ${\epsilon_{\pm}}=\big{(}\Delta_{0}\pm\sqrt{\Omega_{0}^{2}+\Delta_{0}^{2}}\big{)}/2$, for the upper and lower branches, respectively. Their respective (real) spin-$\uparrow$ amplitudes $v$ and $u$ satisfy $u^{2}=\big{(}1+\Delta_{0}/\sqrt{\Omega_{0}^{2}+\Delta_{0}^{2}}\big{)}/2$ and $u^{2}+v^{2}=1$.
@@
$\hat{H}_{\uparrow}=\sum_{\mathbf{k}}(\epsilon_{\mathbf{k}}^{{\vphantom{\dagger}}d}+\nu_{0}^{\vphantom{\dagger}})\hat{d}^{\dagger}_{\mathbf{k}}\hat{d}^{\vphantom{\dagger}}_{\mathbf{k}}+g\sum_{\mathbf{k}\mathbf{q}}\left(\hat{d}^{\dagger}_{\mathbf{q}}\hat{c}^{\vphantom{\dagger}}_{\mathbf{q}-\mathbf{k}\uparrow}\hat{b}^{\vphantom{\dagger}}_{\mathbf{k}}+\hat{b}^{\dagger}_{\mathbf{k}}\hat{c}^{\dagger}_{\mathbf{q}-\mathbf{k}\uparrow}\hat{d}^{\vphantom{\dagger}}_{\mathbf{q}}\right),$ (4)
corresponding to a two-channel model Timmermans et al. (1999) in which two atoms can form a closed-channel dimer $\hat{d}^{\dagger}_{\mathbf{k}}$ with kinetic energy $\epsilon_{\mathbf{k}}^{d}=k^{2}/2(m+M)$ and detuning $\nu_{0}$ from the open channel. This model has the exact low-energy scattering amplitude $f(k)=-1/(a^{-1}+R^{*}k^{2}+ik)$, where the scattering length $a$ and range parameter $R^{*}$ replace the bare model parameters via $m_{r}/2\pi a=-\nu_{0}/g^{2}+\sum_{\mathbf{k}}1/(\epsilon^{b}_{\mathbf{k}}+\epsilon^{c}_{\mathbf{k}})$ and $R^{*}=\pi/m_{r}^{2}g^{2}$, where $m_{r}=mM/(m+M)$ is the reduced mass. The introduction of the closed channel is useful, as it provides a UV cutoff that regulates the Efimov physics, thus introducing an effective three-body parameter. Indeed, for the specific case of a 133Cs-6Li mixture, the deepest bound Efimov trimer unbinds at the scattering length $a_{-}=-1.04R^{*}$ and has energy $E_{0}=-0.242/mR^{*2}$ at unitarity. However, as explicitly shown in the Supplemental Material (SM) SM , our results in the following do not sensitively depend on the choice of UV cutoff  [^1] Our model aims to capture the essential physics of Rabi-coupled mass-imbalanced mixtures. For direct comparison with Cs-Li experiments, one may need to employ more complicated models in order to capture overlapping Fano-Feshbach resonances Repp et al. (2013); Tung et al. (2013), such as models involving three interaction channels Johansen et al. (2017); Yudkin and Khaykovich (2021); Li et al. (2022). or on whether intraspecies scattering is included Ulmanis et al. (2016a).
@@
where $a_{\rm{eff}}$ diverges at $a=a_{\rm{c}}$, as expected. Note that this differs from the resonances induced by sinusoidally driving the scattering length Smith (2015), since there is no imaginary part to $a_{\rm{eff}}$. There also exists an effective scattering length for the upper Rabi-split state; however this is always complex and does not feature resonances Mulkerin et al. (2024); Bleu et al. (2025).
@@
where $E^{\pm}_{\mathbf{k}_{1}\mathbf{k}_{2}}=\epsilon_{\mathbf{k}_{1}}^{b}+\epsilon_{\mathbf{k}_{2}}^{b}+E_{-\mathbf{k}_{1}-\mathbf{k}_{2}}^{\pm}$ and $\gamma_{\mathbf{k}}$ is a three-body vertex function. Similar to the standard three-body problem, we find that the solutions of Eq. (8) yield a spectrum with infinitely many three-body bound states. As shown in Fig. 1(b), these are centered around the new resonance and become unbound at the shifted continuum. In order to directly compare with the usual Efimov scenario, we plot the Rabi-coupled spectrum in terms of the effective scattering length, Eq. (7), and measure energy with respect to the shifted continuum, $\delta E=E-{\epsilon_{-}}$. We see that the spectrum resembles the undriven case, with some modifications for the tightest bound states due to the specific drive parameters chosen, which positions the shifted continuum ${\epsilon_{-}}$ between the lowest- and second-lowest-energy trimers, effectively shifting these states differently. In particular, we find that the excited states satisfy the usual scaling relations for the critical scattering lengths at which the Efimov trimers unbind, $a_{\mathrm{eff},-}^{(n)}=\lambda a_{\mathrm{eff},-}^{(n-1)}$, and for their energies at unitarity, $\delta E^{(n)}=\lambda^{-2}\delta E^{(n-1)}$.
[…]
@@
with $s_{0}\simeq W(1)\sqrt{M/2m_{r}}$ and $k_{\Omega}=\sqrt{2m_{r}\Omega_{0}}$. For a given Efimov trimer, the ratio of binding energies with and without the Rabi drive is given by $\exp\left(2[\delta(R\to\infty)-\delta(R_{0})]/s_{0}\right)$ where $R_{0}$ is a short-range scale SM . Applying this analysis directly yields the purple line in Fig. 2(a), which all trimers with $|E_{0}^{(n)}|\ll\Omega_{0}$ will follow. This matches our exact numerics very well, with small deviations arising from the use of the BO approximation. We have thus demonstrated that the Rabi-dressed three-body parameter is a log-periodic function of the Rabi coupling, and hence that the three-body parameter is tunable via an external drive. The range of possible values is further increased by considering also the role of the detuning SM .
Equation (10) also naturally allows us to consider the modification of inelastic losses due to the Rabi coupling by taking $\delta(R_{0})\to\delta(R_{0})+i\eta$ where $\eta$ is the inelasticity parameter of the uncoupled system. We find that, like the energy, the Rabi-coupled inelasticity parameter is log-periodic with increasing $\Omega_{0}$, and that there are only small oscillations around the bare parameter SM . Therefore, we expect that the Rabi drive would only minimally affect the collisional losses.
@@
Discussion.— Experimentally, one can probe the Efimov spectrum under a continuous Rabi drive by starting in the non-interacting state at $\Delta_{0}/\Omega_{0}\ll-1$ and then adiabatically varying the detuning $\Delta_{0}$ at fixed coupling $\Omega_{0}$, as in Fig. 3. The Efimov states will then show up as loss resonances whenever a bound trimer or tetramer state enters the Rabi-shifted continuum, similarly to how Efimov physics is typically probed using magnetic fields Kraemer et al. (2006). However, since the driving field is typically locked to an atomic frequency standard, this method of exploring Efimov states can potentially yield clearer signatures of highly excited states than in magnetic-field-based experiments. In order to address all the excited Efimov trimers up to the first in a CsLi mixture, i.e., to have $\Omega_{0}$ as large as $|E_{0}^{(1)}|$, we estimate that we need $\Omega_{0}\lesssim 2\pi\times 100\mathrm{kHz}$, which has recently been achieved in Ref. Vivanco et al. (2025). We therefore expect that the modified three-body parameter predicted here can be observed in current or near-future experiments on strongly mass-imbalanced quantum gases.
While our work has focused on the behavior in the steady-state regime, there is also the prospect of using a strong Rabi drive to explore Efimov physics in the time domain, e.g., via Rabi oscillations. This could naturally be implemented for the Bose polaron scenario, where a small number of light impurities are immersed in a gas of heavy bosons. Thus, our few-body results also have potential ramifications for many-body systems such as Bose-Fermi mixtures.
@@ Acknowledgements.
Data Availability.—
The data that support the findings of this article are openly available in the Figshare repository at Zulli et al. (2025).
@@
|$\ket{\Psi_{3}}$ $=\left(\frac{1}{2}\sum_{\mathbf{k}_{1}\mathbf{k}_{2}\sigma}\alpha^{\vphantom{\dagger}}_{\mathbf{k}_{1}\mathbf{k}_{2}\sigma}\hat{c}^{\dagger}_{-\mathbf{k}_{1}-\mathbf{k}_{2}\sigma}\hat{b}^{\dagger}_{\mathbf{k}_{1}}\hat{b}^{\dagger}_{\mathbf{k}_{2}}+\sum_{\mathbf{k}}\gamma^{\vphantom{\dagger}}_{\mathbf{k}}\hat{d}^{\dagger}_{-\mathbf{k}}\hat{b}^{\dagger}_{\mathbf{k}}\right)\ket{0}\,,$ (S1b)
@@
where ${\epsilon_{\pm}}=\big{(}\Delta_{0}\pm\sqrt{\Omega_{0}^{2}+\Delta_{0}^{2}}\big{)}/2$, and the real amplitudes $v$ and $u$ satisfy $u^{2}=\big{(}1+\Delta_{0}/\sqrt{\Omega_{0}^{2}+\Delta_{0}^{2}}\big{)}/2$ and $u^{2}+v^{2}=1$. Finally, replacing the bare interaction parameters with the physical scattering parameters and performing the sum over momenta, we arrive at the following equation:
@@
To look at the effective scattering parameters in the presence of Rabi coupling, we transform from the spin basis to the Rabi-dressed basis and consider the T matrix for the lower Rabi-dressed state:
@@
where we now have an effective scattering length, $a_{\rm{eff}}$, and a modified range parameter $R^{*}_{\rm{eff}}$. This dressed scattering expansion is valid when we have $|\delta E|<\Omega_{0}$. These scattering parameters are both functions of the Rabi-drive parameters and take the form:
@@
where $1/a_{\rm{c}}=v^{2}\sqrt{2m_{r}({\epsilon_{+}}-{\epsilon_{-}})}-2R^{*}{\epsilon_{-}}m_{r}$ is the (inverse) critical scattering length presented in Eq. (6) in the main text. Here, we see that $a_{\rm{eff}}\to\infty$ when $1/a\to 1/a_{\rm{c}}$, as expected, and that the range parameter is modified due to the coupling. In particular, we note that even when $R^{*}\to 0$ we have $R^{*}_{\text{eff}}\neq 0$ in the presence of a Rabi coupling. However, as we discuss below, this Rabi-induced range parameter does not set a new three-body parameter, since this is a purely low-energy property, whereas the Efimov spectrum also relies on the behavior at short distances (high energies).
@@
In the main text, we used a two-channel model which naturally involves a finite effective range that impacts the two-body physics as well as providing an ultraviolet (UV) cutoff for the Efimov spectrum. To show that our key results are model independent, here we instead present all figures using an explicit UV cut-off that only affects the three- and four-body physics Yoshida et al. (2018a, b). To this end, we take the coupling $g$ to infinity while keeping the scattering length $a$ fixed in such a way that $R^{*}=0$. In this limit, the two-body physics coincides with that of a single-channel model. To still provide a three-body parameter for the Efimov spectrum, we follow Ref. Yoshida et al. (2018a) and introduce a UV cut-off $\Lambda$ on the momentum-exchange sums on the right-hand side of the few-body equations (8) and (9) of the main text. The introduction of such a cut-off is equivalent to considering an explicit three-body repulsion Bedaque et al. (1999). To distinguish the two models, we will denote the model with a three-body cut-off as the $\Lambda$ model, as opposed to the two-channel model considered in the main text, which we term the $R^{*}$ model. The deepest bound Efimov cluster for the 133Cs -6Li mixture now enters the scattering continuum at $a_{-}=179.578/\Lambda$.
Figure S1: (a) The Efimov spectra calculated within the $\Lambda$ model with (solid) and without (dashed) Rabi coupling. The two sets of axes represent the uncoupled case (left, top) and the Rabi-driven case (bottom, right), where the latter is expressed in units of the effective scattering length (see Eq. (7) of the main text) and the shifted continuum ($\delta E=E-{\epsilon_{-}}$). The parameters are chosen to correspond to a 133Cs -6Li mixture as in the main text, with $ma_{-}^{2}\Omega_{0}=2$ and $\Delta_{0}/\Omega_{0}=-1$. The uncoupled Efimov spectrum (dashed) is centered around (0,0)(0,0)( 0 , 0 ), while the coupled spectrum (solid) is shifted, and centered instead at $(a_{\text{crit}},{\epsilon_{-}})$ (see arrow). The two-body bound state (purple), three-body trimers (blue) and four-body tetramers (pink) are visible in both spectra. (b) The few-body Rabi-coupled Efimov spectrum as a function of detuning, calculated within the $\Lambda$ model. The parameters correspond to a 133Cs -6Li mixture as in the main text, with $ma_{-}^{2}\Omega_{0}=1$ and $a=|a_{-}|$. Here, we have shaded the scattering continuum (green), the continuum above the dimer (purple) and the continuum above the ground state trimer (blue). The two-body bound state (purple), trimers (blue) and tetramers (pink) all enter the continuum at the lower Rabi-coupled state, with energy ${\epsilon_{-}}$ (dashed line). Inset: The Efimov spectrum centered around the critical detuning ${\Delta_{\text{c}}}$ as a function of $\delta\Delta=\Delta_{0}-{\Delta_{\text{c}}}$.